The diameter of steel rods manufactured on two different extrusion machines is being investigated. Two random samples of sizes n1 = 10 and n2 = 9 are selected, and the sample means and sample variances are 8.73 and 0.35 respectively for sample 1 and 8.68 and 0.40 respectively for sample 2.

Find the lower bound of a 95% two-sided CI on σ1/σ2

Use at least two decimal digits
Answer #1

we have given :- n, -10, 8.73 ts,= 0.35 su for sample Il 9, na = = 8.68 + se =0.40 & for sample 2 =) on +) find 95% CI lower> lower bound 오 F Fi-dla, chei.nirl) 2 Sa here da 1-0.95 = 0.05 .. F 0.975 1 18,9 1 -0.05 0.05 8, 9 0.2295 (use statistical Fo 0.1757 0.2295 = 10.35)2 10.4012 * 0.419178 10.42 95% two sided cI lower bound of on olą =10.42 9

## lower bound of 95 % two sied CI on σ 1 / σ 2 = 0.42